A department store holds a year-end clearance sale that includes a 15.6% discount on cosmetics. Find the sale price of a bottle of perfume if its original price was $47.7.
$

Answer by stanbon(75887)

A department store holds a year-end clearance sale that includes a 15.6% discount on cosmetics. Find the sale price of a bottle of perfume if its original price was $47.7
-----
sale price = 47.7-0.156*47.7 = 0.844*47.7 = $40.26
==========================
Cheers,
Stan H.
============